OpenStudy (anonymous):

5a + 7 = b solve for (a)

OpenStudy (ajprincess):

1.Subtract 7 both sides. 2.Divide by 5 both sides

OpenStudy (anonymous):

5a + 7 = b -7 -7 5a = b-7 /5 /5 a=b/5 -7/5

OpenStudy (anonymous):

so

OpenStudy (anonymous):

\[\frac{ b-7 }{ 5 }\]

OpenStudy (ajprincess):

yup:)

OpenStudy (anonymous):

correct

OpenStudy (anonymous):

Thannks :D